OpenStudy (anonymous):

add and simplify if possible: 49a/7a-49+6a/49a-343

jimthompson5910 (jim_thompson5910):

\[\large \frac{49a}{7a-49}+\frac{6a}{49a-343}\] \[\large \frac{49a}{7(a-7)}+\frac{6a}{49(a-7)}\] \[\large \frac{7*49a}{7*7(a-7)}+\frac{6a}{49(a-7)}\] \[\large \frac{343a}{49(a-7)}+\frac{6a}{49(a-7)}\] \[\large \frac{343a+6a}{49(a-7)}\] \[\large \frac{349a}{49(a-7)}\] \[\large \frac{349a}{49a-343}\] So \[\large \frac{49a}{7a-49}+\frac{6a}{49a-343}=\frac{349a}{49a-343}\]

OpenStudy (anonymous):

thank you, I was killing myself trying to factorize I knew I did not need too.